Shift summary for DAY, Aug. 20, 1999

Angela Bellavance (bellavan@rice.edu)
Fri, 20 Aug 1999 15:33:37 -0500 (CDT)

Shift Summary for day shift, 20/Aug/99

On Shift: Taku (looking curiously like Sydney sometimes), Angela

Hours of beam delivered at all intensities: 5.0833
Hours of beam delivered at nominal intensity: 4.8667
Hours of E832 data (special runs not included): 3.0167
Hours of special runs (laser scans, muon runs...): 0.4, but aborted

Accounting of lost time:
Mainframe 6 controller dropping out
KFRENDs
Pipeline error
Tapes full

========================================================================
Runs ended this shift:
Run 14277 - 23.6M events - BEAM_832
Run 14278 - 1.7M events - BEAM_832
Run 14279 - 2.3M events - BEAM_832
Run 14280-1 - 0 events - Laser scans that didn't init
Run 14282 - 0 events - BEAM_832 (NO good data)
Run 14283 - 0 events - Laser scan initing killed, beam coming.
Run 14284 - 0 events - BEAM_832 (NO good data)
Run 14285 - 0 events - AZLAT testing (nothing written out)
Run 14286 - long LASER SCAN (ABORTED)
Run 14287 - ???? events - BEAM_832 no beam yet but in progress

Tapes brought to FCC this shift: QKC291 - QKC293
QKK291 - QKK293
QKS291 - QKS293
QKB102

========================================================================
Accesses during shift: Julie and Taku to replace mainframe 6 controller.

========================================================================
Overview of shift:
E832 datataking with several problems. Some beam losses due to
machine studies and problems. Shift summary hell.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14277 7:03 Take over shift with smooth running. Taku
dilligently goes to work checking all systems.
324-329 KTeVAna errors (KFREND), trigger in error state,
353 busy, buffer full, more KTeVAna errors.
8:19 Shift crew discover OWL's shift summary still
sitting in an online window. Not sure it was
sent, so resent it.
339-357 Mainframe 6 drops out of CsI. Restarted.
361 KTeVAna errors (KFREND).
389 KTeVAna errors (KFREND).
14277 399 9:05 Tapes fill up, run ended.
14278 1-15 9:12 New run started, but pipeline errors. Also
did a KFREND.
19 Notice online plots missing. Run FIX_DISP_ALL.
14278 40 10:06 MCR calls and says they are doing beam studies.
Beam goes away, stop run.
10:10 Beam unexpectedly returns, and keeps coming.
We call MCR. They _meant_ to say beam _might_
go away.
14279 1 10:22 Do a drt fresh start and begin new run with a
KFREND.
10:25 Taku leaves to take tapes to be vaulted (and to
return his wife's car keys).
19-41 Beam aborts and no beam, but part of some spills
are ok. ~35% beam live time, SEM ~6E12.
11:05 Taku returns.
42-59 11:15 Still some beam aborts and no beam. ~50% live
time, SEM ~3E12.
60 11:42 Beam goes away.
14279 64 11:47 MCR calls, getting experts to look at aborts.
Stop run.
12:00 Beam estimate is 1/2 hour. Laser scan init hangs,
do a drt_fresh_start. Hangs again (kumquat
probs - reset them).
12:21 Beam returns! Give up on laser scan.
14282 1 12:40 Start data-taking run, but mainframe 6 drops out
again. Can't turn back on, page Julie.
12:53 Julie arrives.
12:56 Beam steered away for access.
13:12 Out of access. MCR calls and says He liquifier
down. No beam for a least an hour. What timing!!
13:20 Needed to restore mainframe 8 (messed up in fixing
mainframe 6?).
39 13:26 Run stopped. Set up for laser scan.
13:30 MCR calls - beam returning; kill laser scan init.
14284 1 13:42 Start BEAM_832 run. No beam yet.
13:49 MCR: TeV abort kicker problem, no est.
(It's a conspiracy I say...)
14:00 Gas checklist done. 8 bottles of ArEth and 5
bottles of CO2 delivery in progress.
33 14:24 Run stopped (no spills) so JasonH can work on
AZLAT in slot 12, crate 53.
14285 1 14:32 Init short run to test AZLAT. No problems yet.
14:34 Stop test run.
14286 1 14:43 Start laser scan (for real this time!). Notice
humidity in hall is dropping after Leonard stopped
by and asked about it.
23 15:10 Burns: "Morons! Pathetic morons in my employ
stealing my money!" (i.e. abort laser scan because
beam is back!)
14287 1 15:19 Start data run. KTeVAna errors (KFREND).
2 Beam abort!
3-9 15:50 No beam, hand off to eve shift.

========================================================================
Follow up Notes on Previous Shift

Hey Rick! That "pipeline problem" wasn't a pipeline problem. It was
mainframe 6 dropping out.

========================================================================
Items to Follow up on Next Shift

none

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke